An investment club invested $7,500 into two simple interest accounts for one year. part of the money was invested at 5% and the remainder was placed in a 6% account. The total interest earned was $380. how much was invested into the account?
Answer by Boreal(15235)

x at 5%
7500-x at 6%
interest is .05x+.06(7500-x)
.05x-.06x+450=380
-0.01x=-70
x=$7000 at 5% or $350 interest per year
7500-x=500 @ 6% or $30 interest per year. They add to $380